Factor completely.
9. x(a - b) + 2(a - b)
A) (x + 2)(a - b)
B) (x + 2) + (a - b)
C) (x - 2)(a - b)
D) None of these.

Regroup and factor completely.
mn - nx + my - xy

----------------------------------------------------
I have no idea how to approach these problems please fully explian how to tackel them.




Answer by mszlmb(115)

this is much simpler than it looks, when you have x(a - b) + 2(a - b), for example, one sees the (a-b)'s are the same. so you put them together! now you have x+2 and (a-b), and that's your factoring answer. (x+2)(a-b)
let's check jic!
xa-xb+2a-2b. (i used the foil method)
ok so xa-xb both have n x and
2a-2b both have a 2 so
x(a-b)+2(a-b) is its factor, see, its right! but now we have 2 a-b's we combine them and put the x+2 in ()'s.
(x+2)(a-b) ther4 A is the right answer.
another few examples,
y(2a+c)-v(2a+c)=(y-v)(2a+c)
q(5q+3d)+c(5q+3d)=(q+c)(5q+3d)
get the pattern?